On a range with an elevated tee, a golf enthusiast is practicing. He strikes the golf ball with his club at approximately 20 m/s. The golf ball lands on a point that is 20 m from the edge of the elevated tee. How high is the tee?

1
Expert's answer
2021-03-26T11:37:53-0400
"R=\frac{v^2\sin2\alpha}{g}\\\alpha=0.5\arcsin{\frac{20(9.8)}{20^2}}=15\degree"

"h=\frac{20^2\sin^2(15)}{9.8}=2.7\ m"